OpenStudy (anonymous):

lim goes to -infiniti, [e^(3x) - 2x] / [(1/x) - x]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

use l'hospital's rule

OpenStudy (anonymous):

[3e^(3x) - 2] / [-1/(x^2) - 1] analysing the limit you get - 2 / -1 which is 2

OpenStudy (anonymous):

ok, but if you take 3e^3x to -infiniti it equals infiniti, not zero

OpenStudy (anonymous):

3x^(-infinity)=3/x^(infinity) = 0

OpenStudy (anonymous):

i mean e^(infinity) not x^(infinity)

OpenStudy (anonymous):

wow i was totally missing that the e^infinity drops down to the denominator, whoops! thanks guys